The pH of a weak base 0.10 M is 12.30 What is the Kb of the base and its percent ionization?


What is the H+ of a solution of pyridine, C5H5N, 0.0020 M? Kb = 1.7 x10-9

Calculate the pH of a 0.015 M NaOCl solution and the % hydrolysis of this salt. Ka HOCl = 2.9 x 10-8


One liter of 0.400 M NH3 solution also contains 12.78 g of NH4Cl. How much will the pH of the solution change if 0.075 mol of gaseous HCl is bubbled into it? KbNH3 = 1.8 x 10-5
pH changes from ___________________ to ____________________
1
Expert's answer
2020-08-09T08:29:23-0400

1. pН = 14 – 0.5рКВ + 0.5lgС

0.5рКВ=13.5-12.3

рКВ=2.4

КВ = 3.98x10-3

ionization percent = [OH-]/[Base]= (1012.3-14)/0.1=20%

2. pН = 14 – 0.5рКВ + 0.5lgС = 8.26

[H+]=10-8.26=5.5x10-9

3. [OH] = (Khydrolysis·[NaOCl])1/2

Khydrolysis=Kw/Ka=10-14/2.9x10-8=3.45x10-7

[OH]=7.2x10-5

pH=14-p[OH]=9.86

4. [NH4Cl] = 12.78/53.5=0.24 M

1= 14 - рКb + lg [NH3]/[NH4Cl] = 9.255 + lg(0.4/0.24) = 9.477

2 = 14 - рКb + lg [NH3]/[NH4Cl] = 9.255 + lg(0.4-0.075/0.24+0.075) = 9.268

pH changes from 9.477 to 9.268